# ars548_ros: An ARS 548 RDI radar driver for ROS

### v2

@@ Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long range radar sensors with new RF antenna arrays, which offer digital beam forming. This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1]. Unfortunately, to the best of our knowledge, there are no open source drivers available for Linux systems to enable users to analyze the data acquired by the sensor. In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2). Thus, these data can be stored, represented, and analyzed using the powerful tools offered by ROS. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver including its filtering and representation tools. Besides, we offer a video tutorial to help in its configuration process. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available, so that the user can check the correctness of our driver [2].
@@
Current code version 1.0.0
@@
Software code languages, tools, and services used C++, C.
@@
Current software version 1.0.0
@@ 1 Motivation and significance
One alternative can be the use of thermal infrared cameras for odometry and localization [6, 7]. In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles. Thus, radar sensors are able to penetrate rain, dust and smoke. However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements [8].
@@
In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler-effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications. Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of Doppler effect. This velocity information can be used to improve the odometry estimation in harsh environments [13], which is fundamental for the safe autonomous navigation of robots. Still, their use for 3D navigation is challenging due to the low resolution and narrow vertical Field of View (FoV) of inexpensive models, such as the IWR6843 from Texas Instruments.
More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed. It uses more advanced modulation techniques to better estimate the position and velocity of nearby objects. We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios, as it is a good compromise solution between inexpensive but feature lacking FMCW radar sensors and 2D rotating radars used for automotive purposes. Unfortunately, we did not find any open source driver that would allow us to acquire and process the data acquired from this sensor.
@@
In this paper, we describe the architecture of the ars548_ros package. We highlight its interface to both versions of the Robotic Operating System [14] (ROS), which are becoming the standard for robotic development at least for academic purposes, even though ROS2 is more industrial focused than its predecessor, ROS. To develop ROS and ROS2 interfaces is very convenient as it allows us to use its representation, logging and processing tools. Hereinafter, we will use ROS when we refer to both versions of ROS. We illustrate the good performance of our driver by including data collected from real experiments with measurements of a OS1-32 LiDAR sensor from Ouster as a baseline. Finally, the driver is distributed in the BSD 3-Clause License in the hope that it will be helpful to the community.
@@ 2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects. By using standard messages, the user can take advantage of existing representation and processing tools available in ROS. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual [15]. The messages have been included in Appendix A for the sake of completeness. Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
@@ 2.1 Software architecture
  • ARS548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS548_messages), which are provided to the user via ROS middleware. On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
@@
The full driver source code is publicly available in our GitHub repository [16]. Also, please refer to Appendix B for the additional contents associated with the paper.
@@ 2.2 Software functionalities
This section highlights the main functionalities provided by ARS 548 driver.
@@
With this functionality we provide the user with all the information from ARS548 as emitted by the sensors with our custom messages. Additionally, we provide standard messages which are easier to visualize and process with standard ROS packages.
@@
By generating the PointCloud2 and PoseArray standard ROS messages, the user can easily represent the objects in the standard ROS visualizer (RViz [17]). To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 2a). You can see an example of representation of the radar measurements in Figure 3.
@@ 2.2.3 Object filtering
The ars548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor. The available information is defined in the Object message described in Appendix A. The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details. As an example, see Section 3.2 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
@@ 3 Illustrative examples
The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform [18] (see Figure 4). In particular, we describe here two example scenarios. First, a traffic monitoring scenario, in which the velocity estimation offered by the radar sensor can be of great use to distinguish the road traffic from the background. Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred as Simultaneous Localization and Mapping (SLAM).
@@ 3.1 Experimental setup
Our ARCO robotic platform has been equipped with an ARS 548 RDI sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 4). We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to both sensors via Ethernet, with independent connections. For each experiment, we generated one or more ROSBag files with all the information from both sensors. In particular, with regards to ARS 548 RDI, we saved both our custom topics and the standard ones. The dataset is available at [2].
@@
In this experiment, our ARCO robotic platform remained static in the proximity of a road. The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300m) and its capability of measuring the speed of the detected objects. We use the object filtering feature of our driver (see Figure 2b) obtaining an object point cloud that contains vehicles moving faster than 10 km/h, which are assumed as road traffic (see Figure 5).
Figure 5: Snapshot of Experiment 1. (left) Radar detections on blue, Radar objects on red, Radar object moving faster than 10 km/h on a green sphere and LiDAR measurements on white. (right) Geolocalized radar measurements represented over a map from OpenStreetMap [19]. Red lines represent cumulative radar detections of two different vehicles over the map of the environment.
@@ 4 Impact
This open source driver has been published under the BSD 3-Clause License in the hope of minimizing the development efforts of all users of the ARS 548 RDI sensor who would like to implement ROS-based applications with it. By doing so, the users can easily setup the driver and acquire its data with ROS interfaces and then focus on processing the data as needed for their own experiments, significantly speeding up their research and development process.
@@ 5 Conclusions
Future work will focus on developing new radar odometry approaches, and LiDAR-radar data fusion mechanisms.
@@
All the data used in this paper is openly available at [2].
@@ Acknowledgements
Appendix A . Custom ARS 548 messages
@@
Appendix B . Supplementary data

### v3

@@ Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long-range radar sensors with new RF antenna arrays, which offer digital beam forming. This radar measures independently the distance, speed, and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1]. Unfortunately, to the best of our knowledge, there are no open-source drivers available for Linux systems to enable users to analyze the data acquired by the sensor. In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2). Thus, these data can be stored, represented, and analyzed using the powerful tools offered by ROS. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver including its filtering and representation tools. Besides, we offer a video tutorial to help in its configuration process. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available, so that the user can check the correctness of our driver [2].
@@
Current code version 1.0.1
@@
Software code languages, tools, and services used C++
@@
Current software version 1.0.1
@@ 1 Motivation and significance
One alternative can be the use of thermal infrared cameras for odometry and localization [6, 7], but the problem with cameras is the unknown scale ratio. In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles. Thus, radar sensors are able to penetrate rain, dust and smoke. However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements [8].
@@
In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications. Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of the Doppler effect. This velocity information can be used to improve the odometry estimation in harsh environments [13], which is fundamental for the safe autonomous navigation of robots. Still, their use for 3D navigation is challenging due to the low resolution and narrow vertical Field of View (FoV) of inexpensive models, such as the IWR6843 from Texas Instruments.
More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed. It uses more advanced modulation techniques to better estimate the position and velocity of nearby objects. We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios [23, 24]. This radar is a good compromise solution between inexpensive but feature-lacking FMCW radar sensors and 2D rotating radars used for automotive purposes. Unfortunately, we did not find any open-source driver that would allow us to acquire and process the data acquired from it.
@@
In this paper, we describe the architecture of the ARS 548_ros package. We highlight its interface to both versions of the Robotic Operating System [14] (ROS), which are becoming the standard for robotic development at least for academic purposes, even though ROS2 is more industrially focused than its predecessor, ROS. To develop ROS and ROS2 interfaces is very convenient as it allows us to use their representation, logging, and processing tools. Hereinafter, we will use ROS when we refer to both versions of ROS. We illustrate the good performance of our driver by including data collected from real experiments with measurements of an OS1-32 LiDAR sensor from Ouster as a baseline. Finally, the driver is distributed under the BSD 3-Clause License in the hope that it will be helpful to the community.
@@ 2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information on the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects. By using standard messages, the user can take advantage of existing representation and processing tools available in ROS. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures sent by the sensor as described in the ARS 548 reference manual [15]. The messages have been included in A for the sake of completeness. Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
@@ 2.1 Software architecture
  • ARS 548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS 548_messages), which are provided to the user via ROS middleware. On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor. Finally, the radar_setup utility allows the user to change some parameters of the radar, such as its maximum detection range, its IP address, and many more.
@@
The full driver source code is publicly available in our GitHub repository [16]. Please refer to B for the additional contents associated with the paper.
@@ 2.2 Software functionalities
This section highlights the main functionalities provided by the ARS 548 driver.
@@
With this functionality, we provide the user with all the information from ARS 548 as emitted by the sensors with our custom messages. Additionally, we provide standard messages, which are easier to visualize and process with standard ROS packages.
@@
By generating the PointCloud2 and PoseArray standard ROS messages, the user can easily represent the objects in the standard ROS visualizer (RViz [17]). To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 2a). You can see an example of the representation of the radar measurements in Figure 3.
@@ 2.2.3 Object filtering
The ARS 548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor. The available information is defined in the Object message described in A. The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details. As an example, see Section 3.3 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
2.2.4 Time Stamp Options
We offer two different options to stamp the information received from the ARS 548:
  • Keep the original stamp. Use this option if the ARS 548 device is synchronized with the local PC, which can be achieved by means of the gPTP protocol. To this end, the user PC has to be configured in master mode, so that the sensor can detect it [^1] Please refer to https://linuxptp.sourceforge.net/ for configuring it..
  • Override stamp with local time. In case there is not synchronization between the RADAR sensor and the PC, for example if its Ethernet interface does not support it, we offer an option to override the received timestamp using the local reception time. However, there could be a loss of precision due to incorrect stamping that can be noticeable in high-speed navigation.
2.2.5 Radar configuration
We have included a basic utility that allows the user to change the configuration parameters of the ARS548. The configuration parameters of the device can be divided into the following categories. The reader is referred to the SensorConfiguration.msg file in A to see the detailed list of the configuration parameters:
  1. 1.
    Mounting position. This set of parameters allows the user to configure the mounting position of the device w.r.t. the front axle of the vehicle (lateral, longitudinal and vertical parameters). Besides, the orientation of the device (yaw and pitch angles) and the plug (left or right) can be specified. These parameters are used together with the vehicle parameters in the auto-alignment feature of the ARS 548.
  2. 2.
    Vehicle parameters. They allow the user to specify the main parameters of the car, which are width, length, height and wheel base.
  3. 3.
    Radar configuration. They change the mode of operation of the radar. You can configure its maximum detection distance (from 99 to 1500 meters), choose between three different frequency slots. They can also refer to internal configuration of the Radar, such as its IP address and power saving options, to name a few.
@@ 3 Illustrative examples
The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform [18] (see Figure 4). In particular, we describe here two example scenarios. First, a traffic monitoring scenario, in which the velocity estimation offered by the radar sensor can be of great use to distinguish the road traffic from the background. Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred to as Simultaneous Localization and Mapping (SLAM).
@@ 3.1 Experimental setup
Our ARCO robotic platform has been equipped with an ARS 548 RDI sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 4). We used an MSI GF62 7RE Laptop with 32GB RAM and an 11th generation Intel Core i7 processor. The laptop is connected to both sensors via Ethernet, with independent connections. For each experiment, we generated one or more ROSBag files with all the information from both sensors. In particular, with regards to ARS 548 RDI, we saved both our custom topics and the standard ones. The dataset is available at [2].
3.2 Driver Performance
We have performed a thirty-minute long experiment, in which we used our driver visualized the data from the ARS 548 sensor in real time with RViz. Our goal is to check the usual load and memory usage using the Ubuntu built-in system monitor. During the experiment, the driver process used a mere 2% of the processing power of one core. Furthermore, its memory usage was stable at around 7 MB. Therefore, the driver has a very low impact on the system resources.
@@ 3.2 Driver Performance
In this experiment, our ARCO robotic platform remained static in the proximity of a road. The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300 m) and its capability of measuring the speed of the detected objects. We use the object filtering feature of our driver (see Figure 2b) obtaining an object point cloud that contains vehicles moving faster than 10 km/h, which are assumed to be road traffic (see Figure 5).
Figure 5: Snapshot of Experiment 1. (left) Radar detections in blue, Radar objects in red, Radar object moving faster than 10 km/h in a green sphere and LiDAR measurements in white. (right) Geolocalized radar measurements represented over a map from OpenStreetMap [19]. Red lines represent cumulative radar detections of two different vehicles over the map of the environment.
@@ 4 Impact
This open source driver has been published under the BSD 3-Clause License in the hope of minimizing the development efforts of all users of the ARS 548 RDI sensor who would like to implement ROS-based applications with it. By doing so, the users can easily set up the driver and acquire its data with ROS interfaces and then focus on processing the data as needed for their own experiments, significantly speeding up their research and development process. The code is published in the GitHub platform [16].
@@ 5 Conclusions
Future work will focus on implementing interfaces to allow the ROS modules from the user to provide the ARS 548 RADAR with information, such as estimated velocity and accelerations, to perform its provided auto-alignment function, which is adequate for automotive purposes. Finally, we are developing new radar odometry approaches, and LiDAR-radar data fusion mechanisms.
@@
All the experimental data presented in this paper are openly available at [2].
@@ Acknowledgements
Appendix A Custom ARS 548 messages
@@
SensorConfiguration.msg
1float32 longitudinal #Longitudinal sensor position (AUTOSAR)
2float32 lateral #Lateral sensor position (AUTOSAR)
3float32 vertical #Vertical sensor position (AUTOSAR)
4float32 yaw #Sensor yaw angle (AUTOSAR)
5float32 pitch #Sensor pitch angle (AUTOSAR)
6uint8 plugorientation #Orientation of plug
7float32 length #Vehicle length
8float32 width #Vehicle width
9float32 height #Vehicle height
10float32 wheelbase #Vehicle wheelbase
11uint16 maximumdistance #Maximum detection distance
12uint8 frequencslot #Center frequency (if MaximumDistance < 190 m only Mid can be selected)
13uint8 cycletime #Cycle time
14uint8 timeslot #Cycle offset
15uint8 hcc #Country code
16uint8 powersave_standstill #Power saving in standstill
17uint32 sensoripaddress_0 #Sensor IP address
18uint32 sensoripaddress_1 #Reserved (Configure as 2852025457 or 169.254.116.113)
19uint8 newsensormounting #Flag if new sensor mounting position shall be configured
20uint8 newvehicleparameters #Flag if new vehicle parameters position shall be configured
21uint8 newradarparameters #Flag if new radar parameter shall be configured
22uint8 newnetworkconfiguration # Flag if new IP address shall be configured
Appendix B Supplementary data